OpenStudy (anonymous):

May some one explain to me how to solve a system of equations graphically, graph them, and check? Using "m" and "b"? Here is an example: x+y=6 y=2x-6

OpenStudy (anonymous):

the best way to graph is to put the equation in the [x/a]+[y/b]=1 form rather than y=mx+c form..

OpenStudy (anonymous):

m will decide the slope of your line.. b will decide where ur line cuts the y-axis Draw the two lines Wherever they intersect is the pt of solution

OpenStudy (anonymous):

y=-x+6=>m=-1,b=6 y=2x-6=> m=2 and b=6 the point of intersection is (4,2)

OpenStudy (anonymous):

x+y=6 x/6+y/6=1 u get two pts (0,6),(6,0)...join the line...

OpenStudy (anonymous):

oops,b=-6 for the second equation ...

OpenStudy (anonymous):

For eg y = -x + 6 Draw line with slope -1 and cutting y-axis at (0,6) y=2x - 6 Draw line with slope two cutting y-axis at (0, -6) Mark the pt where these two intersect..That is ur soln in form (x,y)

OpenStudy (anonymous):

from second equation 2x-y=6 x/3+[y/-6]=1 pts...(3,0)and (0,-6) join them nd get the line...

OpenStudy (anonymous):

find the intersection of these two graphs...thats the solution

OpenStudy (anonymous):

for drawing lines graphically, the y=mx+c form is usefull only if u substitute x and find correspoding y. having the slope m and b will not help u plot the graph.

OpenStudy (anonymous):

to plot a graph u need points...and slope intercept form is not so useful here... use the form\[\frac xa+\frac yb=1\]here points of intersection (a,0) and (0,b)...that helps u to plot.

OpenStudy (anonymous):

Alright, thank you everyone for your explanations. I will have to slowly look over everything and try to apply it to the problems I am studying. Thanks very much!